# Mapping tuberculosis fatalities by region and age group in South Korea: A dataset for targeted health policy optimization

### v1

@@ Results
We vary $\vec{\eta}$ to minimize $E_{\rm fatalities}(\vec{\eta})$ while the total number $H$ of hospitals is preserved. The optimal hospital density $\eta_{s}^{\rm(opt)}$ that minimizes $E_{\rm fatalities}(\vec{\eta})$ can, in principle, be obtained analytically using the Lagrange multiplier method. However, when age groups $t$ are taken into account in Eq. (6), a closed-form solution is not available. Therefore, we perform a numerical optimization to determine the optimized hospital configuration vector $\vec{\eta}^{\rm(opt)}=\left(\eta_{1}^{\rm(opt)},\eta_{2}^{\rm(opt)},\cdots,\eta_{s}^{\rm(opt)},\cdots\right)$ using a zero-temperature Monte Carlo approach. A randomly selected pair $(u,v)$ of districts attempts to exchange a small amount of hospital allocation, denoted by $\Delta H$, through the updates $H_{u}\to H_{u}-\Delta H$ and $H_{v}\to H_{v}+\Delta H$. If this relocation results in a decrease in the total fatalities $E_{\rm fatalities}$, the new configuration is accepted. This process is repeated until $E_{\rm fatalities}$ converges to a minimum value, denoted by $E^{\rm(min)}_{\rm fatalities}$. The optimized density configuration is defined as the one that minimizes the objective function, i.e.,
@@
Here we use $\Delta H=0.01$, which is sufficiently small to ensure the convergence of $E^{\rm(min)}_{\rm fatalities}$ with respect to $\Delta H$.
@@
To understand the effect of accounting for age-group distributions on the optimization process, we compare the age-considered optimization results to the baseline case studied in the previous work [11], which uses the objective function in Eq. (1). We first examine the relative change in hospital density compared to the original empirical density $\eta$, specifically ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ and ${\eta^{\rm(opt)}_{\times}}/{\eta}$, which are obtained by minimizing the objective function $E_{\rm fatalities}$ in Eq. (6) for $\eta^{\rm(opt)}_{\rm age}$ and $E(\vec{\eta})$ in Eq. (1) for $\eta^{\rm(opt)}_{\times}$, respectively. Here, the subscript $\times$ represents the optimal hospital density obtained without accounting for age-group differences. Since elderly individuals account for a substantial portion of TB patients and fatalities, we analyze the optimization results with respect to the areal patient density of the oldest age group ($t=5$, ages 80 and above), defined as $\rho_{s,5}=N_{s,5}/A_{s}$. The relative changes ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ and ${\eta^{\rm(opt)}_{\times}}/{\eta}$ are nearly identical [see the color map in Fig. 4(a) and the y-axis scales in Figs. 4(b) and 4(c)].
@@
A higher density of elderly patients alone does not necessarily lead to a greater hospital allocation in the age-considered optimization [Fig. 4(b)], indicating that absolute patient density is not a sufficient predictor of post-optimization hospital gain. In line with Ref. [11], which showed that hospital gain or loss is governed by the characteristic patient density, we therefore examine the optimization outcomes from the perspective of the rescaled density $\rho_{s,5}/\tilde{\eta}_{s,5}$. We find that the rescaled patient density plays a dual role. First, the relative change in hospital density ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ exhibits a clear dependence on ${\rho_{s,5}}/{\tilde{\eta}_{s,5}}$, indicating that this quantity continues to govern whether a district gains or loses hospitals after optimization [Fig. S1 in the Supplemental Material]. Second, the difference between the age-considered and age-agnostic optimizations, ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$, is also systematically controlled by the same rescaled density [Fig. 4(c)]. In this sense, the rescaled patient density ${\rho_{s,5}}/{\tilde{\eta}_{s,5}}$ governs not only the direction of hospital redistribution after optimization, but also amplifies the difference in hospital gains between the age-considered and age-agnostic cases.
The characteristic scale $\tilde{\eta}_{s,t}$ defined in Eq. (5) is strongly correlated with the age-aggregated scale $\tilde{\eta}_{s}$, since the empirical fatality rate $\phi_{s,t}$ is constrained by the overall rate $\phi_{s}$. This correlation explains why the behavior of the hospital density difference also mirrors a similar dependence on the age-aggregated ratio ${\rho}/{\tilde{\eta}}$ [see Fig. S2 in the Supplemental Material].
Figure 4: (a) District-level values of the difference ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$ across South Korea. The inset shows a magnified view of Seoul. Uncolored districts indicate no data satisfying the criteria. (b-c) Comparison of the difference ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$ for $\rho_{s,5}$ (b) and $\rho_{s,5}/\tilde{\eta}_{s,5}$ (c). No distinct pattern observed along $\rho_{s,5}$, while a steady, roughly linear increase is seen along $\rho_{s,5}/\tilde{\eta}_{s,5}$ despite the log-scaled x-axis. (d) The minimum fatalities $E^{\rm(min)}_{\rm{fatalities}}$ as a function of the age weighting parameter $a$ in 2022 is shown representatively, with $a_{\rm min}=-1.4$ and $a_{\rm max}=0.3$. The purple hexagon and green circle represent the cases of $a=a_{\rm min}$ and $a=0$, respectively. The inset figure shows the minimized number of fatalities for each age group $t$ as a function of $a$. (e) The yearly trend of the deviation of $E^{\rm(min)}_{\rm{fatalities}}$ from the original deaths $D$ when $a=a_{\rm min}$ and $a=0$. A bigger value indicates a greater reduction in fatalities compared to the original deaths $D$. Across all analyzed years, the cases with $a=a_{\rm min}$ achieve a better reduction in fatalities than those of $a=0$.
The effect of age scaling in total fatalities
The age distribution of TB patients may vary with time depending on societal factors, such as the baby boom generation or population aging. To account for such scenarios, we introduce a simple linear weight of the form $at+b$ applied to the patient counts in Eq. (6). Incorporating this weighting, the objective function for optimization is redefined as follows:
$E_{\rm{fatalities}}(\vec{\eta};a)=\sum_{s,t}E_{s,t}=\sum_{s,t}(at+b)N_{s,t}\exp\left(-{\eta_{s}}/{\tilde{\eta}_{s,t}}\right),$ (8)
which allows us to rewrite the optimal fatalities $E^{\rm(min)}_{\rm{fatalities}}$ in Eq. (7) as a function of $a$ as $E^{\rm(min)}_{\rm{fatalities}}\equiv E(\vec{\eta}^{\rm(opt)};a)$. We will show later that $b$ is not a free parameter but given by a factor of $a$. The positive value of $a$ assigns greater weight to older age groups, reflecting an aging society, whereas a negative value emphasizes younger age groups, corresponding to a youthful society scenario. When $a=0$, the age-weighted objective function $E_{\rm fatalities}(\vec{\eta};a=0)$ in Eq. (8) reduces to the original, unweighted formulation in Eq. (6). The slope $a$ in the age-weighting function is constrained to lie within a bounded range, $a_{\min}<a<a_{\max}$, in order to avoid unphysical conditions such as negative weighted fatalities for any age group. The bounds are defined as follows:
$a_{\min}$ $=\min[a\in\mathbb{R}\mid at+b\geq 0\text{ for all }t\in\{1,2,3,4,5\}],$
$a_{\max}$ $=\max[a\in\mathbb{R}\mid at+b\geq 0\text{ for all }t\in\{1,2,3,4,5\}],$
where $\mathbb{R}$ is the set of real numbers. The optimized fatalities $E^{\rm(min)}_{\rm{fatalities}}$ can be obtained as a function of $a$ for a given year, and the feasible range of $a$ is approximately $1.4\lesssim a\lesssim 0.3$ over the nine-year period, as shown in Fig. S3 in the Supplemental Material.
Since the optimization process begins from the empirical number of total fatalities, computed as $D=\sum_{s,t}D_{s,t}$ and listed in Table 2, the parameter $b$ in the age-weighting function is chosen to ensure that $E_{\rm{fatalities}}$ evaluated at the empirical $\vec{\eta}$ in Eq. (8) equals the empirical value $D$ for a given $a$. That is, $\sum_{s,t}(at+b)D_{s,t}=D$ and this condition determines $b$ as a function of $a$:
$b(a)=1-a\sum\limits_{s,t}t\,\frac{D_{s,t}}{D}.$ (9)
All quantities on the right-hand side of Eq. (9), except for $a$, are taken from the original empirical values prior to optimization. The characteristic hospital density is not altered by the age scaling because the fatality rate $\phi_{s,t}={D_{s,t}}/{N_{s,t}}$ entering Eq. (5) remains unchanged.
The behavior of the minimum fatalities $E^{\rm(min)}_{\rm{fatalities}}$ as a function of the age-weighting parameter $a$ is shown in Fig. 4(d), using the year 2022 as a representative case. The curve exhibits a slightly concave shape with a peak near $a=0$, which is a typical pattern observed across all years. The respective fatalities $E_{t}^{\rm(min)}$ for each age group tt—leading to $E_{{\rm fatalities}}^{\rm(min)}=\sum_{t=1}^{5}{E_{t}^{\rm(min)}}$—show different behaviors as functions of $a$. The fatalities $E^{\rm(min)}_{5}$ of the oldest group decrease very rapidly with decreasing $a$ for $a<0$, while those of the other groups increase with decreasing $a$, as seen in the inset of Fig. 4(d). The behavior of $E^{\rm(min)}_{5}$ in the oldest group governs the behavior of $E_{{\rm fatalities}}^{\rm(min)}$. The minimum value is $E^{\rm(min)}_{\rm{fatalities}}=731.80$ (approximately a 17% decrease from the empirical total $D=886$) at $a=-1.4\simeq a_{\rm min}$, whereas the corresponding value at $a=0$ is $767.54$ (about a 13% decrease). This implies that the reduction in the elderly patients in a youthful society ($a<0$) would decrease the total fatalities significantly. The tendency $E^{\rm(min)}_{\rm{fatalities}}(a=0)>E^{\rm(min)}_{\rm{fatalities}}(a\simeq a_{\rm min})$ is consistently observed across all periods [Fig. 4(e)]. The curve $E^{\rm(min)}_{\rm{fatalities}}$ as a function of $a$ for all periods is displayed in Fig. S3 in the Supplemental Material.
Figure 5: (a) Relationship between relative changes in optimal hospital allocation $\Delta\eta^{\rm(opt)}_{s}$ versus fatality $\Delta\epsilon^{\rm(opt)}_{s}$ across districts under the $a\simeq a_{\mathrm{min}}$ case compared to the baseline case ($a=0$). (b) A distribution of fatality change $\Delta\epsilon^{\rm(opt)}_{s}$ for age group $t$. $\Delta\epsilon^{\rm(opt)}_{s}$ for $t=5$ (ages 80 and above) concentrated in the negative range, unlike other age groups.
To explore how demographic weighting influences the reallocation of hospitals and its consequences at the district level, we compare two representative scenarios: the empirical baseline with $a=0$ and a hypothetical youthful society with $a\simeq a_{\rm min}$. In both cases, the total number of hospitals remains constant, but the optimal allocation patterns differ due to the application of age-dependent weights in the objective function. For each district $s$, we define the relative changes in hospital allocation and fatalities as follows:
$\Delta\eta_{s}^{\rm(opt)}\equiv\frac{\eta_{s}^{\rm(opt)}(a\simeq a_{\rm min})}{\eta_{s}}-\frac{\eta_{s}^{\rm(opt)}(a=0)}{\eta_{s}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \Delta\epsilon_{s}^{\rm(opt)}\equiv\frac{E_{s}^{\rm(opt)}(a\simeq a_{\rm min})}{D_{s}}-\frac{E_{s}^{\rm(opt)}(a=0)}{D_{s}}.$
These quantities represent, respectively, the relative changes in hospital allocation and total TB fatalities for each district when moving from the baseline to the youthful weighting scenario. Here, the case $a\simeq a_{\rm min}$ is taken as a representative youthful society, while the opposite extreme case $a\simeq a_{\rm max}$, corresponding to an aging society, is analyzed separately in Fig. S4 in the Supplemental Material.
Figure 5(a) shows a scatter plot of $\Delta\eta_{s}^{\rm(opt)}$ versus $\Delta\epsilon_{s}^{\rm(opt)}$ for all districts. In youthful society, some districts lose hospital resources and exhibit reductions in fatalities, whereas others gain resources and show increases in fatalities. To clarify the origin of these different behaviors, we decompose the total fatality change into age-group-specific contributions, $\Delta\epsilon_{t}^{\rm(opt)}$ for $t=1,\dots,5$ [Fig. 5(b)]. This decomposition reveals a strong asymmetry across age groups. The contribution from the oldest group ($t=5$, ages 80 and above) is consistently negative across districts, whereas younger age groups typically show small positive contributions. Consequently, the net fatality change in a district is determined by the balance between the reduction in the oldest group and the cumulative increase among younger groups.
This composition helps explain the outcomes observed under age-weighted optimization. In the youthful weighting scenario, some districts receive more hospitals than in the baseline case and still exhibit an increase in total fatalities. This is not because fatalities rise among the elderly, but because increases in younger age groups outweigh the reduction achieved in the oldest group along with increased hospital resources compared with the baseline scenario. The opposite case is understood in the similar manner. These observations indicate that demographic weighting modifies not only the spatial allocation of hospitals but also how allocation changes translate into fatality outcomes. The comparison between age-weighted and baseline optimizations further shows that demographic weighting can substantially alter spatial allocation patterns under fixed infrastructural constraints (as encoded in $\tilde{\eta}_{s,t}$). Rather than responding uniformly to hospital gains or losses, district-level outcomes reflect a competition between reductions in the oldest age group and cumulative increases across younger groups. This sensitivity to age composition emphasizes the importance of interpreting optimization results through age-resolved contributions rather than aggregate fatality metrics alone, and motivates further discussion on demographic effects in healthcare planning.
Discussion
We present a high-resolution dataset that incorporates age-specific tuberculosis (TB) fatality information and detailed hospital distribution in South Korea. The dataset is constructed by integrating age-stratified ‘si’ and ‘do’ level spatial data with non-age-stratified ‘si’, ‘gun’, and ‘gu’ level spatial data, enabling optimization analyses that reflect both demographic vulnerability and spatial variation at the district scale. A key strength of the dataset is that it supports weighted optimization of healthcare infrastructure through an explicit age-weighting scheme. By introducing an age-weighting parameter into the objective function, we systematically examine how different age-weighting scenarios influence the spatial configuration of hospital resources. While different scenarios can yield similar total numbers of minimized fatalities, the resulting hospital distributions may differ substantially depending on the underlying age structure, highlighting the importance of demographic heterogeneity in spatial healthcare allocation.
In the context of TB, a disease that remains a major public health challenge in South Korea, our analysis demonstrates that age structure plays a critical role in determining optimal spatial allocation patterns. We find that increasing hospital resources in a region does not necessarily lead to proportional reductions in fatalities unless age-dependent risk profiles are taken into account. In some regions, reductions in hospital allocation can still coincide with decreased fatalities, reflecting shifts in age burden or pre-existing differences in characteristic hospital density. These results suggest that spatial optimization strategies that ignore demographic composition may overlook important sources of heterogeneity, particularly in aging societies where regional disparities in age structure and fatality risk are pronounced.
Several limitations of this study should be stated. Age-disaggregated TB data at the district level are reconstructed from province-level distributions, which implicitly assumes homogeneous age composition within each province and may obscure finer-scale demographic variation. In addition, hospital accessibility is represented through areal hospital density and does not explicitly account for differences in capacity, quality of care, or transportation infrastructure. As a result, the optimized configurations should be interpreted as relative allocation patterns rather than as direct prescriptions for policy implementation.
Despite these limitations, the framework provides a consistent basis for comparative or exploratory studies of demographic effects in spatial healthcare allocation. Beyond TB, the weighted formulation of the objective function allows the same approach to be applied to other health-related settings, provided that appropriate data are available. Because the objective function is explicitly formulated as a weighted sum over demographic groups, the same framework can be applied to different diseases or risk profiles by modifying the weighting scheme or the definition of fatality risk, provided that appropriate data are available. In this sense, the dataset can serve as a basis for comparative or exploratory studies rather than as a prescriptive tool for policy design.
@@ Discussion
The code used for data preprocessing, age-stratified fatality modeling, and hospital optimization is openly available at https://github.com/kwyosu7/Tuberculosis_hospital_distribution_optimization. The repository includes Python scripts for constructing the dataset, implementing optimization routines, and generating the figures used in the manuscript.

### v2

@@ Methods
$\vec{\eta}$ is varied to minimize $E_{\rm fatalities}(\vec{\eta})$ while the total number $H$ of hospitals is preserved. The optimal hospital density $\eta_{s}^{\rm(opt)}$ that minimizes $E_{\rm fatalities}(\vec{\eta})$ can, in principle, be obtained analytically using the Lagrange multiplier method. However, when age groups $t$ are taken into account in Eq. (6), a closed-form solution is not available. Therefore, a numerical optimization could be performed to determine the optimized hospital configuration vector $\vec{\eta}^{\rm(opt)}=\left(\eta_{1}^{\rm(opt)},\eta_{2}^{\rm(opt)},\cdots,\eta_{s}^{\rm(opt)},\cdots\right)$ using a zero-temperature Monte Carlo approach. A randomly selected pair $(u,v)$ of districts attempts to exchange a small amount of hospital allocation, denoted by $\Delta H$, through the updates $H_{u}\to H_{u}-\Delta H$ and $H_{v}\to H_{v}+\Delta H$. If this relocation results in a decrease in the total fatalities $E_{\rm fatalities}$, the new configuration is accepted. This process is repeated until $E_{\rm fatalities}$ converges to a minimum value, denoted by $E^{\rm(min)}_{\rm fatalities}$. The optimized density configuration is defined as the one that minimizes the objective function, i.e.,
@@
Here $\Delta H=0.01$, which is sufficiently small to ensure the convergence of $E^{\rm(min)}_{\rm fatalities}$ with respect to $\Delta H$.
@@
To visualize the effect of accounting for age-group distributions on the optimization process, the age-considered optimization results are compared to the baseline case studied in the previous work [11], which uses the objective function in Eq. (3). One can first examine the relative change in hospital density compared to the original empirical density $\eta$, specifically ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ and ${\eta^{\rm(opt)}_{\times}}/{\eta}$, which are obtained by minimizing the objective function $E_{\rm fatalities}$ in Eq. (6) for $\eta^{\rm(opt)}_{\rm age}$ and $E(\vec{\eta})$ in Eq. (3) for $\eta^{\rm(opt)}_{\times}$, respectively. Here, the subscript $\times$ represents the optimal hospital density obtained without accounting for age-group differences. Since elderly individuals account for a substantial portion of TB patients and fatalities, one can analyze the optimization results with respect to the areal patient density of the oldest age group ($t=5$, ages 80 and above), defined as $\rho_{s,5}=N_{s,5}/A_{s}$. The relative changes ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ and ${\eta^{\rm(opt)}_{\times}}/{\eta}$ are nearly identical [see the color map in Fig. 5(a) and the y-axis scales in Figs. 5(b) and 5(c)].
@@
A higher density of elderly patients alone does not necessarily lead to a greater hospital allocation in the age-considered optimization [Fig. 5(b)], indicating that absolute patient density is not a sufficient predictor of post-optimization hospital gain. In line with Ref. [11], which showed that the characteristic patient density governs hospital gain or loss, the optimization outcomes from the perspective of the rescaled density $\rho_{s,5}/\tilde{\eta}_{s,5}$ are examined. We find that the rescaled patient density plays a dual role. First, the relative change in hospital density ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ exhibits a clear dependence on ${\rho_{s,5}}/{\tilde{\eta}_{s,5}}$, indicating that this quantity continues to govern whether a district gains or loses hospitals after optimization [Fig. S1 in the Supplemental Material]. Second, the difference between the age-considered and age-agnostic optimizations, ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$, is also systematically controlled by the same rescaled density [Fig. 5(c)]. In this sense, the rescaled patient density ${\rho_{s,5}}/{\tilde{\eta}_{s,5}}$ governs not only the direction of hospital redistribution after optimization, but also amplifies the difference in hospital gains between the age-considered and age-agnostic cases. The characteristic scale $\tilde{\eta}_{s,t}$ defined in Eq. (5) is strongly correlated with the age-aggregated scale $\tilde{\eta}_{s}$, since the empirical fatality rate $\phi_{s,t}$ is constrained by the overall rate $\phi_{s}$. This correlation explains why the behavior of the hospital density difference also mirrors a similar dependence on the age-aggregated ratio ${\rho}/{\tilde{\eta}}$ [see Fig. S2 in the Supplemental Material].
@@
The code used for data preprocessing, age-stratified data reconstruction, and figure generation is openly available at a github repository: https://github.com/kwyosu7/Tuberculosis_hospital_distribution_optimization. The repository includes Python scripts for constructing the dataset and generating the figures used in the manuscript.